# Time-Series Foundation Model for Value-at-Risk

### v3

@@ 5.1 Actual over Expected Ratio
where $i$ denotes the model in the selected row, whereas $j$ denotes the model in the selected column. Further, the formatting in the Tables are as follows: *, ** and *** denote whether the $t$-test of equal predictive accuracy is rejected at the 5%, 2.5% and 1% level of significance. Table 3 reports the results to $t$-tests. We can observe that the fine-tuned models, in particular, FT1 and FT21, perform remarkably well across all quantile levels, depicted by failure to reject the null hypothesis in the first and second rows of each block. Further, we observe that FT1 outperforms GAS model in forecasting VaR(1%), VaR(5%) and VaR(10%) at 97.5%, 95% and 99% level of significance respectively, and performs atleast as good as GAS model for forecasting Var(2.5%) (refer rows corresponding to GAS model). In addition, we observe that selecting a short horizon as prediction length performs better in predicting VaR. What’s more, the performance of pre-trained models is not so exciting compared to their fine-tuned counterparts. These findings show the potential of time-series foundation models in forecasting value at risk.
@@
Table 3: We report the out-of-sample difference of $|1-AE|$ of each model in the selected column from that of the benchmark in the selected row for Value at Risk at different levels. That is, we report the mean, across the stocks, of $|1-AE|_{i}-|1-AE|_{j}$ where $|1-AE|_{i}$ is the value from model in ith row and $j$ is the value from the model in jth column. We adopt the following formatting: *, ** and *** denotes whether the $t$-test of equal predictive accuracy is rejected at the 5%, 2.5% and 1% level of significance respectively. The hypothesis being tested is $H_{0}:|1-AE|_{i}=|1-AE|_{j}$ against a one-sided alternative $H_{1}:|1-AE|_{i}>|1-AE|_{j}$.
@@
Tables 4, 5 and 6 present the results of these tests, namely UC, CC and DQ respectively, for various confidence levels. The results are also depicted via bar plots in Figures 1-3. For each $VaR_{\alpha}$, we present for how many assets we failed to reject the null hypothesis of correct model performance at various confidence levels. The higher the count better the model. From these tables, it is evident that the TimesFM shows promising results; in particular, the fine-tuned TimesFM model with one prediction length (FT1) either comfortably beats or performs fairly in comparison to the benchmark models across all the three tests. More specifically, for the UC test, the fine-tuned TimesFM models with various prediction lengths and the historical model showed the highest number of assets where the null hypothesis was not rejected, indicating that these models had the most consistent unconditional coverage. What’s more remarkable is that the simple rolling method (historical) outperforms the GAS model in UC test. Again, in conditional coverage tests, the FT model with 1 prediction length performed significantly better than all other models. Finally, the performance of FT1 in DQ test is comparable to the benchmark models if not better. In conclusion, the FT1 model demonstrates strong performance, providing robust competition to the GAS model across multiple tests, while TimesFM models with longer prediction lengths show significant weaknesses, particularly in capturing the conditional and dynamic aspects of the VaR process.
@@
Table 4: This table shows the number of assets (out of 92) for which we failed to reject the null hypothesis in the Unconditional Coverage (UC) test at various Value at Risk (VaR) levels. A higher number indicates better model performance, as it reflects a closer alignment between observed and expected VaR violations for each specified model. Results are reported for four VaR confidence levels (1%, 2.5%, 5%, and 10%) across different models. The best values are highlighted using bold and italicized for the second-best in each row.
@@
Table 5: This table shows the number of assets (out of 92) for which we failed to reject the null hypothesis in the Conditional Coverage (CC) test at various Value at Risk (VaR) levels. A higher number indicates better model performance, as it reflects a closer alignment between observed and expected VaR violations for each specified model. Results are reported for four VaR confidence levels (1%, 2.5%, 5%, and 10%) across different models. The best-performing values are highlighted in bold, while the second-best values are italicized.
@@
Table 6: This table presents the number of assets (out of 92) that passed the Dynamic Quantile (DQ) test at various Value at Risk (VaR) levels across different models and significance levels. A higher count indicates better model performance. The table reports results for four VaR confidence levels (1%, 2.5%, 5%, and 10%). The best-performing values are bolded, and the second-best values are italicized within each row to highlight relative model effectiveness.
@@ 5.2 Quantile Loss
Table 8: We report the out-of-sample total quantile loss of each model in the selected column relative to the benchmark in the selected row for Value at Risk (1%). Each number is a mean of such pairwise relative total quantile losses for each stock. we adopt the following formatting: *, ** and *** denotes whether the Diebold–Mariano (DM) test of equal predictive accuracy is rejected more than 50% of the assets at the 5%, 2.5% and 1% level of significance respectively, across individual tests for each asset. The hypothesis being tested is $H_{0}:\text{total quantile loss}_{i}=\text{total quantile loss}_{j}$ against a one-sided alternative $H_{1}:\text{total quantile loss}_{i}>\text{total quantile loss}_{j}$, where $i$ denotes the model in the selected row, whereas $j$ denotes the model in the selected column.
@@
The cross-sectional mean total quantile score for the fine-tuned model TimesFM, FT1, is below one compared to other TimesFM models, such as FT21 and FT63. However, the results from the DM test do not show a preference for any model over the others. Further, we observe that the mean quantile score decreases as the prediction horizon increases, which suggests that TimesFM is better suited for short-horizon forecasts. The cross-sectional mean total quantile score for the TimesFM fine-tuned models is consistently below one, except for FT63, which scores 1.012 when TimesFM models are evaluated against the historical method as a benchmark. Moreover, the Diebold–Mariano test is rejected at the 5% or 10% significance level for FT1 and FT63 models, indicating that these models perform superior to the historical simulation method. Perhaps the most notable outcome is the competitive performance of the fine-tuned TimesFM models, particularly FT1, against well-established benchmark models such as GAS or Garch-EDF, which are well known for their VaR estimation. Note that the mean quantile score for the GAS model is 1.005 when compared to FT1 as a benchmark. The conclusions remain valid when we compare the performance of these models in forecasting VaR at 2.5% levels. Interestingly, the GAS model is excellent and challenging to beat in forecasting VaR.
We now focus on VaR estimation at 5% and 10% levels, although not very important from the regulator’s perspective, as these capture more common and less severe fluctuations in asset returns and, thus, are less likely to cause significant losses. Nonetheless, we report the comparison. It shows that the TimesFM model fails to beat the historical simulation method at almost all significant levels. What is more surprising compared with extant literature is that the GAS model fails to outperform the Historical Simulation method. A possible reconciliation of this is that the results naturally depend on the data period, split between training and test set, and finally, the results in the literature are usually compared for the stock indices, in contrast to our experiments on multiple stocks constituting the index. Observations for VaR (5%) remain valid in Table for VaR (10%). An appealing but not surprising observation that can be made from results for VaR (10%) is that fine-tuning improves the model performance over the pre-trained model.
@@ 6 Conclusion and discussion
Determining whether pre-trained foundation time-series models can outperform traditional econometric methods is a key question, as their success would mean that anyone with minimal statistical and mathematical knowledge could achieve relatively good VaR estimates. In this paper, we evaluated the performance of Google’s foundation model for time series, TimesFM, in estimating Value-at-Risk (VaR) for the S&P 100 index and its 91 constituents. We initially used the pre-trained TimesFM model without any task-specific training that yielded suboptimal results. However, fine-tuning the model with domain-specific data led to significant statistical improvements: In terms of Actual vs. Expected violations (AE), fine-tuned TimesFM outperforms all econometric benchmarks, including the GAS model, which is widely regarded as the leading method in the recent econometric literature (Patton et al. 2019). Traditional models like GARCH (with different residual distributions) and empirical quantile estimates fall notably behind. These results remain consistent across different VaR levels, underscoring the robustness of TimesFM. Additionally, we assessed performance using the Quantile Score Loss Function, where TimesFM showed results comparable to the best econometric models.
@@
As the results favor the foundation AI model, financial time-series modeling could shift toward data-driven approaches, partially reducing the need for mathematical modeling. Such a shift would have significant implications for both academia and the financial industry, as foundation models can be utilized with relatively minimal methodological expertise in econometrics. However, the adoption of foundation models is accompanied by notable challenges. A key challenge with these models is their “black-box” nature, which can make it difficult to understand how they arrive at their predictions. This is particularly concerning because we still have only a partial understanding of how these models operate, where they might fail, and what they are truly capable of, mainly due to their complex, emergent behaviors. While time series foundation models show promise in adapting to out-of-distribution scenarios, similar to how humans use “common sense” reasoning, they are also prone to hallucinations and may produce decisions that appear reasonable but are incorrect (Chakraborty et al. 2024). This lack of transparency is especially problematic in high-stakes financial settings, where understanding the reasoning behind predictions is essential for effective risk management and decision-making. Moreover, it can create hurdles for regulatory acceptance, as regulators typically require clear and explainable models to meet compliance standards. Therefore, improving the interpretability of foundation models is critical to align with regulatory expectations and support their wider use in practice.
@@
Overall, our results—which are highly generalizable, having been conducted across over 90 assets with multiple settings—favor the pre-trained foundation model, highlighting its potential to enhance Value-at-Risk (VaR) estimation by offering a completely new alternative to traditional econometric methods. We anticipate future research addressing various econometric and financial challenges, such as portfolio optimization and volatility forecasting, while also critically examining the risks associated with applying foundation models to econometric problems.
{APPENDICES}
@@
Figure 1: This figure shows the bar plots of number of assets (out of 92) for which we failed to reject the null hypothesis in the Unconditional Coverage (UC) test at various Value at Risk (VaR) levels. A higher number indicates better model performance, as it reflects a closer alignment between observed and expected VaR violations for each specified model. Results are reported for four VaR confidence levels (1%, 2.5%, 5%, and 10%) across different models.
@@
Figure 2: This figure shows the bar plot of the number of assets (out of 92) for which we failed to reject the null hypothesis in the Conditional Coverage (CC) test at various Value at Risk (VaR) levels. A higher number indicates better model performance, as it reflects a closer alignment between observed and expected VaR violations for each specified model. Results are reported for four VaR confidence levels (1%, 2.5%, 5%, and 10%) across different models.
@@
Figure 3: This figure shows the bar plot of the number of assets (out of 92) for which we failed to reject the null hypothesis in the Dynamic Quantile (DQ) test at various Value at Risk (VaR) levels. A higher number indicates better model performance, as it reflects a closer alignment between observed and expected VaR violations for each specified model. Results are reported for four VaR confidence levels (1%, 2.5%, 5%, and 10%) across different models.

### v4

@@ 5.1 Actual over Expected Ratio
where $i$ denotes the model in the selected row, whereas $j$ denotes the model in the selected column. Further, the formatting in the Tables are as follows: *, ** and *** denote whether the $t$-test of equal predictive accuracy is rejected at the 5%, 2.5% and 1% level of significance. Table 3 reports the results to $t$-tests. We can observe that the fine-tuned models, in particular, FT1 and FT21, perform remarkably well across all quantile levels, depicted by failure to reject the null hypothesis in the first and second rows of each block. Further, we observe that FT1 outperforms GAS model in forecasting VaR(1%), VaR(5%) and VaR(10%) at 97.5%, 95% and 99% level of significance respectively, and performs atleast as good as GAS model for forecasting Var(2.5%) (refer rows corresponding to GAS model). In addition, we observe that selecting a short horizon as prediction length performs better in predicting VaR. What’s more, the performance of pre-trained models is not so exciting compared to their fine-tuned counterparts. These findings show the potential of time-series foundation models in forecasting value-at-risk.
@@
Table 3: We report the out-of-sample difference of $|1-AE|$ of each model in the selected column from that of the benchmark in the selected row for value-at-risk at different levels. That is, we report the mean, across the stocks, of $|1-AE|_{i}-|1-AE|_{j}$ where $|1-AE|_{i}$ is the value from model in ith row and $j$ is the value from the model in jth column. We adopt the following formatting: *, ** and *** denote whether the $t$-test of equal predictive accuracy is rejected at the 5%, 2.5% and 1% level of significance, respectively. The hypothesis being tested is $H_{0}:|1-AE|_{i}=|1-AE|_{j}$ against a one-sided alternative $H_{1}:|1-AE|_{i}>|1-AE|_{j}$.
Tables 4, 5 and 6 present the results of these tests, namely UC, CC and DQ respectively, for various confidence levels. The results are also depicted via bar plots in Figures 1-3. For each $VaR_{\alpha}$, we present for how many assets we failed to reject the null hypothesis of correct model performance at various confidence levels. The higher the count, the better the model. From these tables, it is evident that the TimesFM shows promising results; in particular, the fine-tuned TimesFM model with one prediction length (FT1) either comfortably beats or performs fairly in comparison to the benchmark models across all the three tests. More specifically, for the UC test, the fine-tuned TimesFM models with various prediction lengths and the historical model showed the highest number of assets where the null hypothesis was not rejected, indicating that these models had the most consistent unconditional coverage. What’s more remarkable is that the simple rolling method (historical) outperforms the GAS model in UC test. Again, in conditional coverage tests, the FT model with 1 prediction length performed significantly better than all other models. Finally, the performance of FT1 in DQ test is comparable to the benchmark models if not better. In conclusion, the FT1 model demonstrates strong performance, providing robust competition to the GAS model across multiple tests, while TimesFM models with longer prediction lengths show significant weaknesses, particularly in capturing the conditional and dynamic aspects of the VaR process.
@@
Table 4: This table shows the number of assets (out of 92) for which we failed to reject the null hypothesis in the Unconditional Coverage (UC) test at various Value-at-Risk (VaR) levels. A higher number indicates better model performance, as it reflects a closer alignment between observed and expected VaR violations for each specified model. Results are reported for four VaR confidence levels (1%, 2.5%, 5%, and 10%) across different models. The best values are highlighted using bold and italicized for the second-best in each row.
@@
Table 5: This table shows the number of assets (out of 92) for which we failed to reject the null hypothesis in the Conditional Coverage (CC) test at various value-at-risk (VaR) levels. A higher number indicates better model performance, as it reflects a closer alignment between observed and expected VaR violations for each specified model. Results are reported for four VaR confidence levels (1%, 2.5%, 5%, and 10%) across different models. The best-performing values are highlighted in bold, while the second-best values are italicized.
@@
Table 6: This table presents the number of assets (out of 92) that passed the Dynamic Quantile (DQ) test at various value-at-risk (VaR) levels across different models and significance levels. A higher count indicates better model performance. The table reports results for four VaR confidence levels (1%, 2.5%, 5%, and 10%). The best-performing values are bolded, and the second-best values are italicized within each row to highlight relative model effectiveness.
@@ 5.2 Quantile Loss
Table 8: We report the out-of-sample total quantile loss of each model in the selected column relative to the benchmark in the selected row for value-at-risk (1%). Each number is a mean of such pairwise relative total quantile losses for each stock. we adopt the following formatting: *, ** and *** denotes whether the Diebold–Mariano (DM) test of equal predictive accuracy is rejected for more than 50% of the assets at the 5%, 2.5% and 1% level of significance, respectively, across individual tests for each asset. The hypothesis being tested is $H_{0}:\text{total quantile loss}_{i}=\text{total quantile loss}_{j}$ against a one-sided alternative $H_{1}:\text{total quantile loss}_{i}>\text{total quantile loss}_{j}$, where $i$ denotes the model in the selected row, whereas $j$ denotes the model in the selected column.
The cross-sectional mean total quantile score for the fine-tuned model TimesFM, FT1, is below one compared to other TimesFM models, such as FT21 and FT63. However, the results from the DM test do not show a preference for any model over the others. Further, we observe that the mean quantile score decreases as the prediction horizon increases, which suggests that TimesFM is better suited for short-horizon forecasts. The cross-sectional mean total quantile score for the TimesFM fine-tuned models is consistently below one, except for FT63, which scores 1.012 when TimesFM models are evaluated against the historical method as a benchmark. Moreover, the Diebold–Mariano test is rejected at the 5% or 10% significance level for FT1 and FT63 models, indicating that these models perform superior to the historical simulation method. Perhaps the most notable outcome is the competitive performance of the fine-tuned TimesFM models, particularly FT1, against well-established benchmark models such as GAS or Garch-EDF, which are well known for their VaR forecasting. Note that the mean quantile score for the GAS model is 1.005 when compared to FT1 as a benchmark. The conclusions remain valid when we compare the performance of these models in forecasting VaR at 2.5% levels. Interestingly, the GAS model is excellent and challenging to beat in forecasting VaR.
We now focus on VaR forecasting at 5% and 10% levels, although not very important from the regulator’s perspective, as these capture more common and less severe fluctuations in asset returns and, thus, are less likely to cause significant losses. Nonetheless, we report the comparison. It shows that the TimesFM model fails to beat the historical simulation method at almost all significant levels. What is more surprising compared with extant literature is that the GAS model fails to outperform the Historical Simulation method. A possible reconciliation of this is that the results naturally depend on the data period, split between training and test set, and finally, the results in the literature are usually compared for the stock indices, in contrast to our experiments on multiple stocks constituting the index. Observations for VaR (5%) remain valid in Table for VaR (10%). An appealing but not surprising observation that can be made from results for VaR (10%) is that fine-tuning improves the model performance over the pre-trained model.
@@ 6 Conclusion and discussion
Determining whether pre-trained foundation time-series models can outperform traditional econometric methods is a key question, as their success would mean that anyone with minimal statistical and mathematical knowledge could achieve relatively good VaR forecasts. In this paper, we evaluated the performance of Google’s foundation model for time series, TimesFM, in estimating Value-at-Risk (VaR) for the S&P 100 index and its 91 constituents. We initially used the pre-trained TimesFM model without any task-specific training that yielded suboptimal results. However, fine-tuning the model with domain-specific data led to significant statistical improvements: In terms of Actual vs. Expected violations (AE), fine-tuned TimesFM outperforms all econometric benchmarks, including the GAS model, which is widely regarded as the leading method in the recent econometric literature (Patton et al., 2019). Traditional models like GARCH (with different residual distributions) and empirical quantile estimates fall notably behind. These results remain consistent across different VaR levels, underscoring the robustness of TimesFM. Additionally, we assessed performance using the Quantile Score Loss Function, where TimesFM showed results comparable to the best econometric models.
@@
As the results favor the foundation AI model, financial time-series modeling could shift toward data-driven approaches, partially reducing the need for mathematical modeling. Such a shift would have significant implications for both academia and the financial industry, as foundation models can be utilized with relatively minimal methodological expertise in econometrics. However, the adoption of foundation models is accompanied by notable challenges. A key challenge with these models is their “black-box” nature, which can make it difficult to understand how they arrive at their predictions. This is particularly concerning because we still have only a partial understanding of how these models operate, where they might fail, and what they are truly capable of, mainly due to their complex, emergent behaviors. While time series foundation models show promise in adapting to out-of-distribution scenarios, similar to how humans use “common sense” reasoning, they are also prone to hallucinations and may produce decisions that appear reasonable but are incorrect (Chakraborty et al., 2024). This lack of transparency is especially problematic in high-stakes financial settings, where understanding the reasoning behind predictions is essential for effective risk management and decision-making. Moreover, it can create hurdles for regulatory acceptance, as regulators typically require clear and explainable models to meet compliance standards. Therefore, improving the interpretability of foundation models is critical to align with regulatory expectations and support their wider use in practice.
@@
Overall, our results—which are highly generalizable, having been conducted across over 90 assets with multiple settings—favor the pre-trained foundation model, highlighting its potential to enhance Value-at-Risk (VaR) forecasting by offering a completely new alternative to traditional econometric methods. We anticipate future research addressing various econometric and financial challenges, such as portfolio optimization and volatility forecasting, while also critically examining the risks associated with applying foundation models to econometric problems.
7 Declaration of Generative AI and AI-assisted technologies in the writing process
While preparing this work, the authors used GPT-4 to refine the language and enhance the clarity of the manuscript. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
8 CRediT author statement
Anubha Goel: Data curation; Formal analysis; Investigation; Validation; Visualization; Writing- Original draft preparation. Puneet Pasricha: Formal analysis; Investigation; Methodology; Validation; Visualization; Writing- Original draft preparation. Juho Kanniainen: Conceptualization; Methodology; Project administration; Supervision; Validation; Visualization; Writing- Original draft preparation; Writing- Reviewing and Editing.
9 Acknowledgements
The first author acknowledges support from the European Union’s Horizon Europe programme under the Marie Skłodowska-Curie Actions (Grant Agreement No. 101150609, Project: HiddenTipChains).
@@ 9 Acknowledgements
Figure 1: This figure shows the bar plots of number of assets (out of 92) for which we failed to reject the null hypothesis in the Unconditional Coverage (UC) test at various value-at-risk (VaR) levels. A higher number indicates better model performance, as it reflects a closer alignment between observed and expected VaR violations for each specified model. Results are reported for four VaR confidence levels (1%, 2.5%, 5%, and 10%) across different models.
@@
Figure 2: This figure shows the bar plot of the number of assets (out of 92) for which we failed to reject the null hypothesis in the Conditional Coverage (CC) test at various value-at-risk (VaR) levels. A higher number indicates better model performance, as it reflects a closer alignment between observed and expected VaR violations for each specified model. Results are reported for four VaR confidence levels (1%, 2.5%, 5%, and 10%) across different models.
@@
Figure 3: This figure shows the bar plot of the number of assets (out of 92) for which we failed to reject the null hypothesis in the Dynamic Quantile (DQ) test at various value-at-risk (VaR) levels. A higher number indicates better model performance, as it reflects a closer alignment between observed and expected VaR violations for each specified model. Results are reported for four VaR confidence levels (1%, 2.5%, 5%, and 10%) across different models.